OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. 2x + y = 3 -3x + 4y = 1 (1, 1) (-1, -1) (1, -1) (-1, 1)

OpenStudy (abb0t):

Did you attempt to solve this problem yourself first?

OpenStudy (anonymous):

No I did not.

OpenStudy (anonymous):

y=-2x + 3 -3x + 4(-2x + 3) = 1 -3x + -8x + 12 = 1 -11x = -11 x = 1 so y = 1

OpenStudy (anonymous):

next time try it on your self

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

its not that hard or easy